OpenStudy (anonymous):

I roll a die twice. What is the probability that I will get a bigger number in the second toss?

OpenStudy (anonymous):

@campbell_st @jim_thompson5910

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

take all cases

OpenStudy (anonymous):

i was thinking that lol

OpenStudy (misty1212):

first roll is a one, then second could be 2,3,4,5,6 so \[\frac{1}{6}\times \frac{5}{6}\] for that one

OpenStudy (misty1212):

first roll is a two, second could be 3,4,5,6 so \[\frac{1}{6}\times \frac{4}{6}\] for that one etc

OpenStudy (anonymous):

okay so 5/36 +4/36 +3/36 +2/36+1/36 ?

OpenStudy (anonymous):

so 15/36 ?

OpenStudy (anonymous):

am i right satelitie ?

OpenStudy (anonymous):

i would go with that-, yes

OpenStudy (anonymous):

thanks guys!

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Lucky Louie flipped a coin 8 times. What is the probability that it came up tails every time (all 8 times) ?

OpenStudy (anonymous):

damned unlikely

OpenStudy (anonymous):

last one i promise lol

OpenStudy (anonymous):

\[\frac{1}{2^8}\]

OpenStudy (anonymous):

A single card is drawn from an ordinary deck of 52 cards. Find the probability that the card is a face card (jack, queen or king) OR a heart.